OpenStudy (babyslapmafro):

Please help me determine whether the given differential equation is linear or nonlinear and give the order of the equation. click to see... thanks

OpenStudy (babyslapmafro):

\[x^2dy+(y-xy-xe^x)dx=0\]

OpenStudy (amistre64):

are any y parts have an exponent greater than 1?

OpenStudy (amistre64):

\[x^2dy+(y-xy-xe^x)dx=0\] \[x^2~y'+(y-xy-xe^x)=0\] might be easier to see

OpenStudy (amistre64):

the order relates to the highest derivative; wheres the linear aspect of it relates to the exponents and a few other things like trig functions and what not.